OpenStudy (anonymous):

f(x) = 3x^3+11x^2-67x+21 What are the 3 real zeros of this polynomial ??

myininaya (myininaya):

try find the rational zeros by finding the factors of 21

myininaya (myininaya):

\[f(1)=3+11-67+21=14+21-67=35-67 \neq 0\] so try -1,3,-3,-7,7,-21,21

myininaya (myininaya):

you are wanting to see if one these will give you f=0

OpenStudy (anonymous):

i have 3 and -7 that equal zero but i am having trouble finding the 3rd zero

OpenStudy (anonymous):

the last one is a fraction? i think

myininaya (myininaya):

oh yeah forgot to include some other factors like factors of 21/factors of 3 but that is good that you found two zeros

myininaya (myininaya):

we only need one of those zeros to write this as a factor(quadratic)=0

myininaya (myininaya):

3| 3 11 -67 21 | 9 60 -21 ================ 3 20 -7 | 0

myininaya (myininaya):

\[3x^2+20x-7=0\]

myininaya (myininaya):

\[3x^2+21x-1x-7=0\] \[3x(x+7)-1(x+7)=0\]

myininaya (myininaya):

\[(x+7)(3x-1)=0\]

myininaya (myininaya):

any questions?

OpenStudy (anonymous):

the last complex zero of this polynomial is 1/3 thanks :)

myininaya (myininaya):

that is right! :)

OpenStudy (anonymous):

How do i use the complex zeros for factor F ?

myininaya (myininaya):

you want to factor the polynomial?

myininaya (myininaya):

\[f(x)=(x+7)(3x-1)(x-3)\]

myininaya (myininaya):

if x=a,b,c are the zeros of f, then f=(x-a)(x-b)(x-c)

OpenStudy (anonymous):

i got 3(x-1/3)(x+7)(x-3) thanks :)

myininaya (myininaya):

yes that is fine you can write it like that :)